OpenStudy (anonymous):

Choose the equation that could be used to find two consecutive integers whose sum is 67. n + (n + 2) = 67 n − (n + 1) = 67 n + n = 67 n + (n + 1) = 67

hero (hero):

Which do you believe is the correct choice?

OpenStudy (anonymous):

C

hero (hero):

Well, no, that's not how consecutive integers work.

hero (hero):

If you have two consecutive integers, the second integer is always one more than the first.

hero (hero):

For example: 5 and 6 are consecutive integers. So if 5 = n, then 6 = n + 1

OpenStudy (anonymous):

sooo A?

hero (hero):

Now you're just guessing...I just explained it to you above. There's only one possible correct choice.

hero (hero):

You already know 5 + 6 = 11. If you didn't know what the integers were, you would write n + (n + 1) = 11

OpenStudy (anonymous):

oh my bad D

hero (hero):

Just to let you know, I only like explaining things once.

hero (hero):

So try to think things through before answering